OpenStudy (anonymous):

what two numbers when multiplied = to 2, but the sum of is 0? two numbers multiplied equals to 8, but added together equals to 5 i'm really confused, thanks for the help

ganeshie8 (ganeshie8):

say the required two numbers are \(x\) and \(y\)

ganeshie8 (ganeshie8):

you want the numbers to be multiplied to `2` ? and the sum to equal `0` right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good, setup the equations accordingly

ganeshie8 (ganeshie8):

\(xy = 2\) \(x+y = 0\)

ganeshie8 (ganeshie8):

solve them ^

OpenStudy (anonymous):

um, that's the part where i'm confused at

ganeshie8 (ganeshie8):

my display pic may help u, take a look :)

ganeshie8 (ganeshie8):

eliminate one of the variables first : solve for x from the second equation and plug it in the first equation

ganeshie8 (ganeshie8):

second equation : \(x+ y = 0\) \(x = ?\)

OpenStudy (anonymous):

are the answers decimals?

ganeshie8 (ganeshie8):

idk yet, we need to solve and see

OpenStudy (anonymous):

x=2?

ganeshie8 (ganeshie8):

not quite \(x + y = 0\) \(x = -y\)

ganeshie8 (ganeshie8):

plug this in first equaiton : \(xy = 2\) \((-y)y = 2\) \(-y^2 = 2\) \(y^2 = -2\)

ganeshie8 (ganeshie8):

can the square of a `real ` number be negative ever ?

ganeshie8 (ganeshie8):

nope, so there are `no solutions` in real numbers

OpenStudy (anonymous):

so this problem doesn't have an answer?

ganeshie8 (ganeshie8):

the answer is `no solutions` in real numbers

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

is this ur second question ? two numbers multiplied equals to 8, but added together equals to 5

OpenStudy (anonymous):

yes, it doesn't have a solution either right?

ganeshie8 (ganeshie8):

you cant tell it just by looking... its difficult lol

ganeshie8 (ganeshie8):

setup the equations and work it like we did the earlier problem

ganeshie8 (ganeshie8):

but you're right... there are no solutions for that also

OpenStudy (anonymous):

the only factors of 8 are 1,2,4,8, and neither of those numbers would equal 5 even if we had a negative

OpenStudy (anonymous):

ok, thanks

ganeshie8 (ganeshie8):

*no real solutions : http://www.wolframalpha.com/input/?i=solve+xy+%3D+8%2C+x%2By%3D5

ganeshie8 (ganeshie8):

very good :) thats a nice approach !

OpenStudy (anonymous):

the original problem was find the roots of each quadratic polynomial: x^2-6x-7 (x__)(x___) and this was the way my teacher showed us

OpenStudy (anonymous):

wait, wrong problem *x^2+2*

OpenStudy (anonymous):

well, thanks for the help, and bye!